OpenStudy (anonymous):

assuming x doesn't equal 0 and y doesnt equal 0 what is the quotient of 8x^9y^3+2x^6y^2+8x^3y over 2x^3y please help me!

OpenStudy (shamil98):

Take the gcf out of the numerator and reduce

OpenStudy (anonymous):

ohhh ok thanks!

OpenStudy (anonymous):

wait im still confused

OpenStudy (anonymous):

willing to help?

OpenStudy (anonymous):

I really need help this one and another one!

OpenStudy (campbell_st):

well start by rewriting the fraction \[\frac{8x^9y^3}{2xy^3} +\frac{2x^6y^2}{2xy^3}+\frac{8xy^3}{2xy^3}\]

OpenStudy (anonymous):

ok I get it so far

OpenStudy (campbell_st):

then start with numbers simplify the numbers then use the index law for division \[\frac{x^a}{x^b} = x^{a -b}\] and you'll get the quotient

OpenStudy (anonymous):

would the answer be 4x^27y^3+x^18y^2+4x^9y?

OpenStudy (anonymous):

LaTex is not working at this moment. Refer to the attachment.

OpenStudy (anonymous):

thank you for showing your work robtobey its going to help me when I study! you guys where a big help :) have a really good day!

OpenStudy (campbell_st):

well the numbers are correct... letter are correct but the powers aren't looking at the first fraction its \[\frac{8x^9y^3}{2xy^3} = 4x^{9 -1}y^{3-3} = 4x^8y^0 = 4x^8\] does this make sense

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

I have 2 questions I need help with: 1) what is the slope of a line that passes through the point (-1,1) and is parallel to a like that passes through (4,6) and (-1,-4) 2) Blake has a total of 64 DVD's and CD's. If the number of the CD's is 8 more times the number of DVDs how many DVD's does he have?

OpenStudy (anonymous):

@campbell_st If it was revealed that the Numerator terms are in reverse order, would the solution then be correct? Mathematica reverses expression terms, not to confuse the user, but because that is the way they chose to code the program. They have a function to "make it look right", called TraditionalForm.

OpenStudy (anonymous):

ok... im confused what do you mean by that..i need help on those 2 problems I typed earlier